### v1

# Random close packing fraction of bidisperse discs: Theoretical derivation and exact bounds

Abstract
A long-standing problem has been a theoretical prediction of the densest packing fraction of random packings, $\Phi_{RCP}$, of same-size discs in $d=2$ and spheres in $3$. However, to minimize order, experiments and numerical simulations often use two-size discs. For practical purposes, then, a predictive theory for the packing fraction, $\Phi_{RCP}$, of the densest such bidisperse packings is more useful. A disorder-guaranteeing theory is formulated here to fill this gap, using an approach that led to an exact solution for monodisperse discs in $d=2$ Bl21 . $\Phi_{RCP}$ depends on the sizes ratio, $D$, and concentrations, $p$, of the disc types and the developed theory enables derivation of exact and rigorous upper and lower bounds on $\Phi_{RCP}(p,D)$, as well as an explicit prediction of it.
Amongst the many variants of packing problems, the random close packing (RCP) of monodisperse spheres in three dimensions and discs in two have become canonical problems, relevant to the fields of mathematics, physics, and engineering. Loosely posed, the quest is to predict the densest disordered state that such packings can pack into. Other than useful to technological applications Toetal10 and liquid-solid phase transitions Bernal , a successful theoretical approach to address this problem can pave the way to predicting the densest packing of more general objects. Addressing the RCP problem requires a disorder-ensuring criterion Toetal00 because same-size spheres and discs tend to crystalize into ordered states under most packing processes, which are denser than disordered ones. Many trial-and-error experimental and numerical studies put the highest packing fraction, $\Phi_{RCP}$, inside a narrow range of values. However, this does not rule out a yet-untried process that could generate a denser state. Given that all packing processes occupy an infinite-dimensional parameter space, testing them individually is not feasible. Another difficulty is ensuring disorder. Several disorder criteria have been proposed Toetal00 . However, these are difficult to implement because they often require packing particles first and then testing the level of disorder. An ideal criterion should be integral to an a-priori prediction of $\Phi_{RCP}$. A recent theoretical approach to predict $\Phi_{RCP}$ for packings of monodisperse discs in $d=2$ overcame both difficulties by using the cell order distribution (COD) to replace the infinite process parameter space and resulted in a rigorous exact prediction Bl21 .
While the problem of the monodisperse RCP appears resolved for discs in $d=2$, many simulations and experiments are carried out on packings of two-size, or bidisperse, discs to minimize crystallization ViBo72 ; Bietal86 ; OkOd04 ; Xuetal05 ; Koetal16 ; Br25 . Therefore, predicting $\Phi_{RCP}$ theoretically in such systems is, arguably, more useful. The bidisperse problem is more complex because, in addition to the difficulties of the infinite process parameter space and limiting order, $\Phi_{RCP}$ in such packings is not a number but a function of the disc sizes ratio, $D$, and the concentration of the smaller discs, $p$. Nevertheless, the approach developed for monodisperse packings paves the way to construct this more involved theory. Developing this approach is the main aim here. The following is limited to discs of $\left(1,D_{max}\right)$ in $d=2$, where $D_{max}=\sqrt{3}/(2-\sqrt{3})$ ensures no ‘rattlers’ within enclosures of three large discs.
In the following, the COD is first defined and the advantages of using it are briefly reviewed. An exact upper bound on $\Phi_{RCP}$ is then derived as a function of $p$ and $D$. Next, the disorder criterion, developed in Bl21 , is extended and used to identify the range of $p(D)$, within which packings are assured to be disordered. An exact calculation of $\Phi_{RCP}(p,D)$ follows, based on the identification of the COD of the densest packing. The final result is a derivation of an exact lower bound on $\Phi_{RCP}(p,D)$. A summary of the results and a discussion conclude the paper.
The COD: Without loss of generality, the diameter of the smaller discs is taken to be unity and their concentration $p$. Any theory, aiming to predict $\Phi_{RCP}$, faces the problem of sensitivity to the packing process, which exists in an infinite parameter space. This difficulty is alleviated by using the COD MaBl14 ; Waetal20 , defined as follows. Connecting the centres of discs in contact generates a graph, whose nodes are the disc centers and edges are the lines connecting them. The smallest voids enclosed by the edges are the ’cells’, the number of edges surrounding a cell is its order, and its distribution is the COD,
$P(k)=\sum_{j=3}^{\infty}Q_{j}\delta_{kj}\ ,$ (1)
with $\delta_{kj}$ the Kronecker delta function and $Q_{k}$ the fraction of cells of order $k$ (henceforth, $k$-cells) out of all existing $N_{c}$ cells.
The first advantage of the COD is that it correlates directly with the density - increasing the fraction of low-order cells increases the mean number of contacts per disc, the density, and the packing fraction. Secondly, as any packing process generates a COD, it effectively parameterizes all possible packings. This circumvents the infinite parameter space problem and predictions based on the COD hold for all processes. Thirdly, it can be used to determine directly the packing fraction of any packing of $N$ discs and $N_{c}$ cells, as follows.
$S_{pack}=N_{c}\sum_{k=3}^{\infty}Q_{k}\bar{S}_{k}\equiv N_{c}s_{pack}\ ,$ (2)
where $\bar{S}_{k}$ is the mean area of $k$-cells over all their configurations and $s_{pack}$ is the mean packing area per cell irrespective of its order. To vitiate boundary corrections, which decay as $1/\sqrt{N}$ in $d=2$, only the disc areas contained within the cells are considered for calculating the packing fraction. Recalling that the internal angles of a $k$-cell sum up to $(k-2)\pi$,
$\Phi$ $=\frac{N_{c}\sum_{k=3}^{\infty}Q_{k}\frac{(k-2)\pi}{8}\left[p+(1-p)D^{2}\right]}{N_{c}\sum_{k=3}^{\infty}Q_{k}\bar{S}_{k}}$
$=\frac{\pi\left[p+(1-p)D^{2}\right]\left(\bar{k}-2\right)}{8s_{pack}}\ ,$ (3)
where $\bar{k}=\sum_{k=3}^{\infty}kQ_{k}$ is the mean of the COD. Thus, given $p$ and $D$, the solution of the random packing problem amounts to finding the ‘ideal’ COD that maximizes (3) under the constraint of disorder.
Exact upper bound on $\Phi_{RCP}$: The densest possible packings contain only $3$-cells. While ordered such packings can be generated Fe25 , it is unclear whether disordered ones are topologically possible. Nevertheless, such hypothetical packings provide a rigorous upper bound for $\Phi_{RCP}(p,D)$.
Figure 1: The four possible 33-cell configurations and their areas (shaded).
33-cells come in four configurations, whose areas, shown shaded in Fig. 1, are
$S_{3a}$ $=\frac{\sqrt{3}}{4}\quad;\quad S_{3b}=\frac{\sqrt{D\left(D+2\right)}}{4}$
$S_{3c}$ $=\frac{D\sqrt{1+2D}}{4}\quad;\quad S_{3d}=\frac{\sqrt{3}}{4}D^{2}\ .$ (4)
The areas, which the discs occupy inside these triangular areas, are:
$S^{disc}_{3a}$ $=\frac{\pi}{8}\quad;\quad S^{disc}_{3b}=\frac{\arccos{\frac{1}{D+1}}+D^{2}\arcsin{\frac{1}{D+1}}}{4}$
$S^{disc}_{3c}$ $=\frac{D^{2}\arccos{\frac{D}{D+1}}+\arcsin{\frac{D}{D+1}}}{4}\quad;\quad S^{disc}_{3d}=\frac{\pi D^{2}}{8}\ .$ (5)
It is convenient to parameterize the occurrence probabilities of the four configurations by the auxiliary probability $0\leq u\leq 1$, with v=1uv=1-u:
$\left(p_{3a},\ p_{3b},\ p_{3c},\ p_{3d}\right)=\left(u^{3},\ 3u^{2}v,\ 3uv^{2},\ v^{3}\right)Q_{3}\ .$ (6)
For $D\neq 1$, $u\neq p$ because adjacent $3$-cells are correlated through the sharing of two discs. The exact relation $p(u)$ is derived below. As $N\to\infty$, the total area of this packing is
$\bar{S}_{3}=$ $N_{c}\sum_{i}p_{3i}S_{3i}=\frac{\sqrt{3}N_{c}}{4}\bigg[u^{3}+v^{3}D^{2}+$
$uv\sqrt{3}\left(\sqrt{D(D+2)}u+D\sqrt{2D+1}v\right)\bigg]\ ,$ (7)
of which the discs occupy
$\bar{S}^{disc}_{3}$ $=N_{c}\sum_{i}p_{3i}S^{disc}_{3i}=\frac{\pi N_{c}}{8}\Bigg[\left(u^{3}+v^{3}D^{2}\right)+$
$\frac{6uv}{\pi}\bigg[u\left(\arccos{\frac{1}{D+1}}+D^{2}\arcsin{\frac{1}{D+1}}\right)+$
$v\left(D^{2}\arccos{\frac{D}{D+1}}+\arcsin{\frac{D}{D+1}}\right)\bigg]\Bigg]\ .$ (8)
When $N\to\infty$, $\bar{S}^{disc}_{3}=\pi\left[p+(1-p)D^{2}\right]N/4$ up to corrections of order $\cal{O}$$(\sqrt{N})$. Equating this expression with (8) and noting that $N_{c}=2N$ in such packings yields a relation $p(u,D)$:
$p=u^{3}+\frac{6uv}{\pi}\left(u\arccos{\frac{1}{D+1}}+v\arcsin{\frac{D}{D+1}}\right)\ ,$ (9)
up to corrections of order $\cal{O}$$(1/\sqrt{N})$. This relation is plotted in the supplemental material SM for ten values of $1.5\leq D\leq 6.0$. Using this relation, a plot of $\Phi_{3}(p,D)=\bar{S}^{disc}_{3}(p,D)/\bar{S}_{3}(p,D)$ is shown in Fig. 2. The line connecting the maxima of the curves provides the absolute maximum of $\Phi_{3}$, $\Phi_{3max}(p_{max},D)$, for any $D$ and the concentration $p_{max}$ at which it is attained. Each of these curves is an upper bound, $\Phi_{3}(p,D)\geq\Phi_{RCP}(p,D)$.
$\Phi_{3}$ also provides the actual packing fraction of ordered $3$-cell bidisperse disc packing, each of which corresponding to a specific combination of $p$, $D$, and $Q_{3}$ Fe25 .
Figure 2: The packing fraction, $\Phi_{3}(p,D)$, for $1.0\leq D\leq 6.0<D_{max}=\sqrt{3}/(2-\sqrt{3})$. The blue line follows the highest packing fraction for each mixture (add ’max’ in its legend). $D=1.o$ recovers the fully trigonal crystal, $\Phi_{3}=\pi/\left(2\sqrt{3}\right)$.
A disorder criterion: Order often appears in the form of large clusters of trigonal lattices of identical $3$-cells, i.e., of clusters comprising only either configuration $a$ or $d$ in Fig. 1. The occurrence probability of such clusters increases with $Q_{3}$ and, close to $p=0$ or $1$, such packings cannot be considered disordered. To identify the allowed range of $p$ within which disorder is assured, a criterion, developed for monodisperse packings, is extended next. It states that, if the probability that a $3$-cell has, on average, only one identical neighboring cell, then the occurrence probability of clusters of $L$ such cells, $P(L)$, decays exponentially with $L$, with $P(L>5)<1.73\times 10^{-4}$ Bl21 . The occurrence probabilities of configurations $a$ and $d$ are, respectively, $u^{3}Q_{3}$ and $v^{3}Q_{3}$ and the probability of a cell of either configuration having more than one identical neighbor is
$R_{a}=u^{3}Q_{3}\left[1-\left(u^{3}Q_{3}\right)^{3}-3\left(1-u^{3}Q_{3}\right)\left(u^{3}Q_{3}\right)^{2}\right]$
$R_{d}=v^{3}Q_{3}\left[1-\left(v^{3}Q_{3}\right)^{3}-3\left(1-v^{3}Q_{3}\right)\left(v^{3}Q_{3}\right)^{2}\right]\ .$ (10)
Imposing that neither probability exceeds $1/3$, $R_{a}<1/3$ and $R_{d}<1/3$, (10), yields the conditions for for $u^{3}Q_{3}$ and $v^{3}Q_{3}=(1-u)^{3}Q_{3}$
$Q_{3}<min\left\{\frac{0.562236...}{u^{3}},\frac{0.562236...}{\left(1-u\right)^{3}},1\right\}\ .$ (11)
These provide the value of $Q_{3max}(u)$. From (11), $Q_{3}<1$ when 0.82535265>u>10.825352645=0.174647350.82535265>u>1-0.825352645=0.17464735. Expressing this condition in terms of $p$ and $D$, using eq. (9), provides the allowed range of $p$ for all $D$, plotted in Fig. 3. In particular, keeping to $0.311749<p<0.82535265$, assures disorder for any value of $D$. This criterion holds for any packing. For example, disorder is assured for the frequently used ratio $D=\sqrt{2}$ when $0.202521367<p<0.85220001$. Another example, is the common choice that the two disc types occupy the same area, $p/[p+(1-p)D^{2}]=0.5$, which corresponds to $p=1/(D^{2}+1)$. Calculating the disorder range for such packings and plotting it also in Fig. 3, shows that this practice runs a higher risk of crystallization for $D\lesssim 3$.
Figure 3: For any choice of $D$, a 3-cells system would be disordered when $u$ is between $u_{min}(D)$ (purple curve) and $u_{max}(D)$ (green curve). The lowest curve extends from $p_{min}(D=1)=0.17464735$ to $p_{min}(D=6.4)=0.311749$ and the upper curve from $p_{max}(D=1)=0.82535265$ to $p_{max}(D=6.4)=0.938588$. It follows that, for any value of $1<D<D_{max}$, such systems would be disordered when $p_{low}\equiv 0.311749<p<0.82535265\equiv p_{high}$. Also shown is the curve corresponding to the common choice of both disc types occupying the same area (blue curve). The sharp drop of the upper bound in these packings, when $D\lesssim 3$, limits the range of $p$, for which disorder is assured.
The densest packing fraction: Anticipating that disordered packings of only $3$-cells may be topologically impossible, the ideal packing should comprise only $3$- and $4$-cells, whose packing fractions are
$\Phi=\frac{Q_{3}\bar{S}^{disc}_{3}+\left(1-Q_{3}\right)\bar{S}^{disc}_{4}}{Q_{3}\bar{S}_{3}+\left(1-Q_{3}\right)\bar{S}_{4}}\ .$ (12)
$\Phi_{RCP}(p,D)$ is achieved when $Q_{3}=Q_{3max}(p,D)$ and, to calculate it, requires the averages $\bar{S}^{disc}_{4}(p,D)$ and $\bar{S}_{4}(p,D)$. These are derived in detail in the supplemental material SM . While $\Phi_{RCP}(p,D)$ can be calculated to any precision, using (12), a closed form for it is very cumbersome and it was calculated numerically for a hundred values of $p$ and ten values of $1\leq D\leq D_{max}$. A typical example for $D=3.0$ is shown in Fig. 4 and the aggregated results for the values of $D$ in Fig. 5. In both these figures, the plots of $\Phi_{RCP}(p,D)$ are flanked by the upper bound, derived above, and the lower bound, derived next.
Figure 4: A typical example of $\Phi_{RCP}$ of bidisperse planar disc packing as a function of $p$ for $D=3.0$. It is bounded above by $\Phi_{3}$ and below by $\Phi_{low}$.
Figure 5: As in Fig. 5 for several size ratios $1.0\leq D\leq 6.0$. $\Phi_{RCP}(p,D)$ are plotted together with their respective upper and lower bounds.
Within a central region of $p$, $\Phi_{RCP}$ coincides with the upper bound for all $D$. This is because $Q_{3max}=1$ there and those packings are disordered even if they could comprise, hypothetically, only $3$-cells. This issue is discussed in more detail below. Significantly, an inspection of the plots reveals that, for every size ratio, the maximal possible value of $\Phi_{RCP}$ coincides with the upper bound. An observation that should aid planning compositions of bidisperse packings.
Exact lower bound on $\Phi_{RCP}$: Focusing on packings of only $3$- and $4$-cells, it should be noted that the packing are disordered for all $1\leq D<D_{max}$ and any value of $p$ when $Q_{3max}=0.562236$.... As $\Phi$ is a monotonically increasing function of $Q_{3max}$, substituting this value in relation (12) provides an exact and rigorous lower bound on $\Phi_{RCP}$, irrespective of $D$ (even when $D\to 1$),
$\Phi_{low}=\frac{0.562236\bar{S}^{disc}_{3}+0.437763\bar{S}^{disc}_{4}}{0.562236\bar{S}_{3}+0.437763\bar{S}_{4}}\ .$ (13)
These bounds are also plotted in Figs. 4 and 5.
To conclude, a theory, based on the cell order distribution (COD), has been developed to solve analytically the long-standing problem of random close packing of bidisperse discs in $d=2$, for size ratios that exclude rattlers in $3$-cells. A disorder-assuring criterion has been derived, which is an extension of the one developed for monodisperse disc packings Bl21 . This criterion allows to determine, for any disc size ratio, $D$, the range of small disc concentration, $p$, within which the packing is disordered. This information is useful for avoiding crystallization when designing experiments and simulations of disordered disc systems. In particular, it was shown that a common choice of bidisperse systems whose two disc types occupy the same area, increases the risk of crystallisation for $D\lesssim 3$. Exact $p$- and $D$-dependent upper and lower bounds on $\Phi_{RCP}$ have been derived, the former corresponding to a disordered packing of only $3$-cells and the latter corresponding to the highest fraction of $3$-cells at which disorder is assured for all $D$. Then the exact value of the random close packing fraction, $\Phi_{RCP}$, was derived as a function of $p$ and $D$. It has been found that, for every size ratio, the highest possible packing fraction always coincides with the upper bound. The packing fraction, at which this occurs, $p_{3max}$, can serve as a guide to future physical and numerical experiments in such systems. The use of the COD obviates the sensitivity to the packing process, which introduces an infinite parameter space that is impossible to explore fully by trial and error. It follows that the results obtained here are valid in general, irrespective of the packing protocol.
In setting the disorder criterion, only trigonal order has been considered. While could be argued that large clusters of some $4$-cells may also represent order, being deformed square lattices, it is shown in the supplemental material SM that a disorder criterion, similar to (11), can also be derived for $Q_{4}$ and that the ideal COD determined above always satisfies it.
These results are useful to guide experiments and simulations, which rely on disordered disc packings. They can also be used to signal formation of order, when measured packing fractions are too high. Nevertheless, it should be emphasized that aiming to derive the densest possible packing fraction, $\Phi_{RCP}$, is independent of whether or not such dense packings can be generated realistically. In particular, it is unclear whether or not packings with $Q_{3max}>0.562236$... are possible topologically. This issue, which is downstream from this paper, remains to be investigated. Interestingly, examining, where possible, existing simulations aiming to generate the densest disordered bidisperse disc packings ViBo72 ; Bietal86 ; OkOd04 ; Xuetal05 ; Koetal16 ; Br25 , in none $Q_{3}$ appears to be higher than this value. This may be one of the reasons that observed packing fractions of bidisperse disc packings are often lower than the theoretical values obtained here. It is possible that planning experiments with the values of $p_{3max}(D)$, identified here, may result in higher packing fractions.
Finally, the method developed here is useful beyond predicting the densest disordered bidisperse disc packings. As has been noted in passing, the method yields the packing fractions of ordered packings that comprise of only $3$-cells Fe25 . Moreover, it can be used to analyze general multidisperse disc packings of any COD. Such analyses require determining the statistics of cell orders higher than $4$ and, while this could be demanding to achieve analytically or in close form, such calculations can be readily done numerically.

### v2

@@
 Abstract
-A long-standing problem has been a theoretical prediction of the densest packing fraction of random packings, $\Phi_{RCP}$, of same-size discs in $d=2$ and spheres in $3$.
+A long-standing problem has been a theoretical prediction of the densest packing fraction of random packings, $\Phi_{RCP}$, of same-size discs in $d=2$ and spheres in 33.
 However, to minimize order, experiments and numerical simulations often use two-size discs.
@@
 Therefore, predicting $\Phi_{RCP}$ theoretically in such systems is, arguably, more useful.
-The bidisperse problem is more complex because, in addition to the difficulties of the infinite process parameter space and limiting order, $\Phi_{RCP}$ in such packings is not a number but a function of the disc sizes ratio, $D$, and the concentration of the smaller discs, $p$.
+The bidisperse problem is more complex because, in addition to the difficulties of the infinite process parameter space and limiting order, $\Phi_{RCP}$ in such packings is not a number but a function of the disc sizes ratio, $D$, and the concentration of the smaller discs, pp.
 Nevertheless, the approach developed for monodisperse packings paves the way to construct this more involved theory.
@@
 A summary of the results and a discussion conclude the paper.
-The COD: Without loss of generality, the diameter of the smaller discs is taken to be unity and their concentration $p$.
+The COD: Without loss of generality, the diameter of the smaller discs is taken to be unity and their concentration pp.
 Any theory, aiming to predict $\Phi_{RCP}$, faces the problem of sensitivity to the packing process, which exists in an infinite parameter space.
@@
 This circumvents the infinite parameter space problem and predictions based on the COD hold for all processes.
-Thirdly, it can be used to determine directly the packing fraction of any packing of $N$ discs and $N_{c}$ cells, as follows.
+Thirdly, it can be used to determine directly the packing fraction of any packing of NN discs and $N_{c}$ cells, as follows.
 $S_{pack}=N_{c}\sum_{k=3}^{\infty}Q_{k}\bar{S}_{k}\equiv N_{c}s_{pack}\ ,$ (2)
@@
 Thus, given $p$ and $D$, the solution of the random packing problem amounts to finding the ‘ideal’ COD that maximizes (3) under the constraint of disorder.
-Exact upper bound on $\Phi_{RCP}$: The densest possible packings contain only $3$-cells.
+Exact upper bound on $\Phi_{RCP}$: The densest possible packings contain only 33-cells.
 While ordered such packings can be generated Fe25 , it is unclear whether disordered ones are topologically possible.
@@
 $\left(p_{3a},\ p_{3b},\ p_{3c},\ p_{3d}\right)=\left(u^{3},\ 3u^{2}v,\ 3uv^{2},\ v^{3}\right)Q_{3}\ .$ (6)
-For $D\neq 1$, $u\neq p$ because adjacent $3$-cells are correlated through the sharing of two discs. The exact relation $p(u)$ is derived below. As $N\to\infty$, the total area of this packing is
+For $D\neq 1$, $u\neq p$ because adjacent 33-cells are correlated through the sharing of two discs. The exact relation $p(u)$ is derived below. As $N\to\infty$, the total area of this packing is
 $\bar{S}_{3}=$ $N_{c}\sum_{i}p_{3i}S_{3i}=\frac{\sqrt{3}N_{c}}{4}\bigg[u^{3}+v^{3}D^{2}+$
@@
 Each of these curves is an upper bound, $\Phi_{3}(p,D)\geq\Phi_{RCP}(p,D)$.
-$\Phi_{3}$ also provides the actual packing fraction of ordered $3$-cell bidisperse disc packing, each of which corresponding to a specific combination of $p$, $D$, and $Q_{3}$ Fe25 .
+$\Phi_{3}$ also provides the actual packing fraction of ordered 33-cell bidisperse disc packing, each of which corresponding to a specific combination of $p$, $D$, and $Q_{3}$ Fe25 .
 Figure 2: The packing fraction, $\Phi_{3}(p,D)$, for $1.0\leq D\leq 6.0<D_{max}=\sqrt{3}/(2-\sqrt{3})$.
-The blue line follows the highest packing fraction for each mixture (add ’max’ in its legend).
+The blue line follows the highest packing fraction for each mixture.
 $D=1.o$ recovers the fully trigonal crystal, $\Phi_{3}=\pi/\left(2\sqrt{3}\right)$.
-A disorder criterion: Order often appears in the form of large clusters of trigonal lattices of identical $3$-cells, i.e., of clusters comprising only either configuration $a$ or $d$ in Fig.
+A disorder criterion: Order often appears in the form of large clusters of trigonal lattices of identical 33-cells, i.e., of clusters comprising only either configuration $a$ or $d$ in Fig.
 1.
-The occurrence probability of such clusters increases with $Q_{3}$ and, close to $p=0$ or $1$, such packings cannot be considered disordered.
+The occurrence probability of such clusters increases with $Q_{3}$ and, close to $p=0$ or 11, such packings cannot be considered disordered.
 To identify the allowed range of $p$ within which disorder is assured, a criterion, developed for monodisperse packings, is extended next.
-It states that, if the probability that a $3$-cell has, on average, only one identical neighboring cell, then the occurrence probability of clusters of $L$ such cells, $P(L)$, decays exponentially with $L$, with $P(L>5)<1.73\times 10^{-4}$ Bl21 .
+It states that, if the probability that a 33-cell has, on average, only one identical neighboring cell, then the occurrence probability of clusters of $L$ such cells, $P(L)$, decays exponentially with $L$, with $P(L>5)<1.73\times 10^{-4}$ Bl21 .
 The occurrence probabilities of configurations $a$ and $d$ are, respectively, $u^{3}Q_{3}$ and $v^{3}Q_{3}$ and the probability of a cell of either configuration having more than one identical neighbor is
@@
 The sharp drop of the upper bound in these packings, when $D\lesssim 3$, limits the range of $p$, for which disorder is assured.
-The densest packing fraction: Anticipating that disordered packings of only $3$-cells may be topologically impossible, the ideal packing should comprise only $3$- and $4$-cells, whose packing fractions are
+The densest packing fraction: Anticipating that disordered packings of only 33-cells may be topologically impossible, the ideal packing should comprise only 33- and 44-cells, whose packing fractions are
 $\Phi=\frac{Q_{3}\bar{S}^{disc}_{3}+\left(1-Q_{3}\right)\bar{S}^{disc}_{4}}{Q_{3}\bar{S}_{3}+\left(1-Q_{3}\right)\bar{S}_{4}}\ .$ (12)
@@
 Within a central region of $p$, $\Phi_{RCP}$ coincides with the upper bound for all $D$.
-This is because $Q_{3max}=1$ there and those packings are disordered even if they could comprise, hypothetically, only $3$-cells.
+This is because $Q_{3max}=1$ there and those packings are disordered even if they could comprise, hypothetically, only 33-cells.
 This issue is discussed in more detail below.
@@
 An observation that should aid planning compositions of bidisperse packings.
-Exact lower bound on $\Phi_{RCP}$: Focusing on packings of only $3$- and $4$-cells, it should be noted that the packing are disordered for all $1\leq D<D_{max}$ and any value of $p$ when $Q_{3max}=0.562236$....
+Exact lower bound on $\Phi_{RCP}$: Focusing on packings of only 33- and 44-cells, it should be noted that the packing are disordered for all $1\leq D<D_{max}$ and any value of $p$ when $Q_{3max}=0.562236$....
 As $\Phi$ is a monotonically increasing function of $Q_{3max}$, substituting this value in relation (12) provides an exact and rigorous lower bound on $\Phi_{RCP}$, irrespective of $D$ (even when $D\to 1$),
@@
 These bounds are also plotted in Figs. 4 and 5.
-To conclude, a theory, based on the cell order distribution (COD), has been developed to solve analytically the long-standing problem of random close packing of bidisperse discs in $d=2$, for size ratios that exclude rattlers in $3$-cells.
+To conclude, a theory, based on the cell order distribution (COD), has been developed to solve analytically the long-standing problem of random close packing of bidisperse discs in $d=2$, for size ratios that exclude rattlers in 33-cells.
 A disorder-assuring criterion has been derived, which is an extension of the one developed for monodisperse disc packings Bl21 .
@@
 In particular, it was shown that a common choice of bidisperse systems whose two disc types occupy the same area, increases the risk of crystallisation for $D\lesssim 3$.
-Exact $p$- and $D$-dependent upper and lower bounds on $\Phi_{RCP}$ have been derived, the former corresponding to a disordered packing of only $3$-cells and the latter corresponding to the highest fraction of $3$-cells at which disorder is assured for all $D$.
+Exact $p$- and $D$-dependent upper and lower bounds on $\Phi_{RCP}$ have been derived, the former corresponding to a disordered packing of only 33-cells and the latter corresponding to the highest fraction of 33-cells at which disorder is assured for all $D$.
 Then the exact value of the random close packing fraction, $\Phi_{RCP}$, was derived as a function of $p$ and $D$.
@@
 In setting the disorder criterion, only trigonal order has been considered.
-While could be argued that large clusters of some $4$-cells may also represent order, being deformed square lattices, it is shown in the supplemental material SM that a disorder criterion, similar to (11), can also be derived for $Q_{4}$ and that the ideal COD determined above always satisfies it.
+While could be argued that large clusters of some 44-cells may also represent order, being deformed square lattices, it is shown in the supplemental material SM that a disorder criterion, similar to (11), can also be derived for $Q_{4}$ and that the ideal COD determined above always satisfies it.
 These results are useful to guide experiments and simulations, which rely on disordered disc packings.
@@
 Finally, the method developed here is useful beyond predicting the densest disordered bidisperse disc packings.
-As has been noted in passing, the method yields the packing fractions of ordered packings that comprise of only $3$-cells Fe25 .
+As has been noted in passing, the method yields the packing fractions of ordered packings that comprise of only 33-cells Fe25 .
 Moreover, it can be used to analyze general multidisperse disc packings of any COD.
-Such analyses require determining the statistics of cell orders higher than $4$ and, while this could be demanding to achieve analytically or in close form, such calculations can be readily done numerically.
+Such analyses require determining the statistics of cell orders higher than 44 and, while this could be demanding to achieve analytically or in close form, such calculations can be readily done numerically.